# Hyperparameter Optimization in Machine Learning

### v2

@@ 7.3 Hypergradients via iterative differentiation
Another approach to computing hypergradients involves explicitly accounting for an iterative training dynamics. Methods based on iterative differentiation further assume that solutions to the learning algorithm problem (7.3) are sought (with approximation) by iterating an optimization dynamics for $T$ steps:
@@
We call $s_{t}=[w^{\mathsf{T}}_{t},v^{\mathsf{T}}_{t}]^{\mathsf{T}}\in\mathds{R}^{d}$ the “state” of the optimizer, which contains the current parameters, $w_{t}$ and a (possibly empty) set of auxiliary variables the $v_{t}$ – e.g. velocities and moments for algorithms such as Adam and RMSProp. This process yields a solution $w_{T}=w_{T}(\lambda)$. The transition maps $\Phi_{t}:(\mathds{R}^{d}\times\mathds{R}^{m}\times\mathscr{D})\rightarrow\mathds{R}^{d}$ are assumed to be differentiable and describe the actual operations associated with a specific optimizer. Typically, $\Phi_{t}$ depends explicitly on the current minibatch $\mathcal{D}_{t}$ (e.g. for stochastic gradient descent) and it internally requires the computation of the inner objective gradients
@@
Note that $\Phi$ depends both explicitly and via $s_{t-1}$, and $\lambda$ might contain parameters related to the optimizer itself such as the learning rate. As an example, the dynamics (7.9) for stochastic gradient descent with momentum are
@@
In a nutshell, iterative differentiation schemes for gradient-based HPO rely on the application of automatic differentiation (AD) to the composite map $\mathscr{M}\circ\Phi_{T}\circ\dots\circ\Phi_{1}$. We briefly survey AD next, and refer the reader to [Griewank08:_evaluat_deriv] for an in-depth account.
@@ 7.3.1 Automatic differentiation
Automatic (or algorithmic) differentiation (AD) [Griewank08:_evaluat_deriv, baydin2017automatic] is the standard tool for computing gradients in deep learning. Suppose we have a generic differentiable function $G:\mathds{R}^{N}\mapsto\mathds{R}^{M}$ with $(\eta_{1},\dots,\eta_{M})=G(\xi_{1},\dots,\xi_{M})$ and we are interested in computing the Jacobian matrix, with entries $\frac{\partial\eta_{i}}{\partial\xi_{j}}$. In AD, we begin by describing $G$ as a computational graph $(V,E)$, which is a directed acyclic graph whose nodes $V$ include inputs $\xi_{j}$, outputs $\eta_{i}$, and intermediate variables that are computed by applying elementary functions to inputs or other intermediate variables, that can be scalars, vectors or tensors. Elementary functions are typically atomic operations such as (tensor) sum, (Hadamard) product, and (elementwise) trascendental functions, whose partial derivatives are readily known. Edges $E$ in the graph describe functional dependencies, i.e., they link intermediate variables to the elementary functions they are fed into. Forward propagation, following a topological sort of the graph, allows us to compute $G(\xi_{1},\dots,\xi_{M})$. Two AD strategies are available, forward and reverse mode, which are based on the definition of local derivatives and a message-passing procedure based on the standard chain rule of calculus.
@@
initialized as $\dot{\xi_{j}}=1$, where $\pi_{k}$ are $v_{k}$’s parents, and iterated following a topological sort of $(V,E)$. In the above equation, if $v_{k}$ and $v_{\ell}$ are vector-valued, $\frac{\partial v_{k}}{\partial v_{\ell}}$ denotes a Jacobian matrix. The desired derivatives can be finally read as $\dot{\eta}_{i}$. The same propagation allows us to compute the $j$-th column of the Jacobian and to obtain the whole Jacobian, propagation needs to be repeated for each input $\xi_{j}$. Assuming elementary functions can be computed in $O(1)$, the total computational cost is $O(NE)$. Forward mode is thus advantageous for functions with few inputs and many outputs. If on the other hand $N=O(E)$, as it happens for example for the objective function of neural networks, forward mode is as inefficient as computing derivatives numerically. It is therefore not surprising that forward mode is never used to train neural network models (although there is a remarkable exception when training recurrent networks on data streams [williams_learning_1989], a setting where reverse mode cannot be applied).
In reverse mode (the strategy used by the well known backpropagation algorithm [rumelhart1986learning]) the local derivative of $v_{k}$, usually called adjoint, is defined as
@@ 7.3.2 Computing hypergradients
We are now ready to apply the principles of AD to the differentiation of the validation error with respect to $\lambda$. A simplified view of the computational graph mapping $\lambda$ into the response $f(\lambda)$ is shown in Figure 7.1, where minibatches $\mathcal{D}_{t}$ and initial weights $w_{0}$ are supplied as given constants [^2] Note that $w_{0}$ could be equally well included in $\lambda$ but here we keep it distinct for clarity.. Here the number of inputs $M$ is the dimension of the hyperparameter vector and there is a scalar ($N=1$) output. Forward and reverse calculations are presented in [franceschi2017forward] and repeated below.
@@
$\nabla f(\lambda)=\frac{\partial\mathscr{M}(w_{T}(\lambda),\mathcal{V})}{\partial\lambda}\dot{s}_{T}$
@@
$\bar{s}_{T}=\frac{\partial\mathscr{M}(w_{T}(\lambda),\mathcal{V})}{\partial\lambda}$
@@ 7.4 Fixed-point methods
Another possible approach that sits in between the iterative and the implicit differentiation schema is given by the so-called fixed-point method. The method is based on the observation that one can rewrite the condition (7.4) as a fixed point equation of an appropriate contractive dynamics
@@
which could be, for instance, gradient descent $\Phi(w,\lambda)=w-\eta\nabla\mathcal{J}$.  [^3] As in the iterative case, the dynamics may be defined on an augmented state, for example when using gradient descent with momentum. Then, the application of the implicit function theorem to (7.15) yields to
@@
and, repeating the procedure described above, Equation (7.7) becomes
@@
Starting from any point $q_{T,0}\in\mathds{R}^{d}$, one can iterate
@@
to find an approximate fixed-point of (7.16) and then plug $q_{T,K}$ into Equation (7.8) to obtain an hypergradient.
This computational scheme emerges in more general cases where $\Phi$ is not necessarily an optimization dynamics: $\Phi$ may represent a stable recurrent neural network [miller2019stable], a graph neural network [scarselli2009graph] or an equilibrium model [bai2019deep]. In these contexts, the fixed-point algorithm is usually known as recurrent backpropagation [almeida1987learning, liao2018reviving].
Warm-starting $q_{T,0}$ with the vector obtained at the previous optimization iteration (of $\lambda$) may speed-up the convergence of (7.17).
@@ 7.5 Use cases
One possibility is to attach a hyperparameter to every weight in a deep network. We have already mentioned the idea proposed by maclaurin2015gradient of regarding the initial values for the weights as hyperparameters. In the same paper, the authors also suggest a closely related opportunity where a regularization penalty is associated to every individual weights. The meta-learning technique presented in [franceschi2018bilevel] can be interpreted as treating all weights in the first layers of a deep network as hyperparameters, and optimize the loss on new tasks with respect to the weights in the classification head.
Another direction that has been studied is to attach hyperparameters to data. maclaurin2015gradient demonstrated an extreme experiments where image pixels themselves are hyperparameters. In a more realistic scenario, hypergradients have been used to tune the hyperparameters of differentiable data augmentation procedures, yielding significantly speedups without performance loss compared to other data augmentation searching methods [lin:2019:online_hyperparameter_learning, hataya:2020:faster_autoaugment_learning]. A different opportunity is to assign a cost weight to each example in a supervised learning setting so that the overall loss for a training set of $n$ examples is $\sum_{i=1}^{n}\lambda^{(i)}\mathcal{J}(h(x^{(i)}),y^{(i)})$. Here we have $n$ hyperparameters $\lambda^{(i)}$. franceschi2017forward show how this idea can be exploited to filter out from the training set examples with noisy (or perhaps poisoned) labels. In this regards, it is interesting to note that (although not directly related to HPO) the problem of building training set attacks on linear models can be formulated in terms of bilevel optimization [mei:2015:using_machine_teaching] and has commonalities with gradient-based HPO in its mathematical structure.
@@
7.6 Probabilistic reparameterization of discrete hyperparameters
The presence of discrete hyperparameters into the response function – and therefore into the training objective or dynamics – prevents a straightforward application of gradient-based HPO techniques. For some such hyperparameters, however, it might make sense to reformulate them as random variable that follow some discrete distribution. For instance, consider a discrete binary choice $\lambda\in\{0,1\}$, e.g. indicating whether to use an $L^{1}$ or $L^{2}$ regularizer in (2.3). Then, one may define a discrete binary random variable $\iota$ as follows:
@@ 7.6 Probabilistic reparameterization of discrete hyperparameters
and reinterpret $\lambda$ as a drawing of $\iota$, where $\mathrm{Ber}(\theta)$ is the Bernoulli distribution with success parameter $\theta$. 444 Similarly, one may reformulate hyperparmeters encoding multiple choices via categorical distributions. Crucially, the distribution parameter $\theta$ is now continuous and one may consider optimizing the response function
@@
instead of $f$, having that $\min_{\theta}g(\theta)=\min_{\lambda}f(\lambda)$. Although (7.19) might look like a circular argument, the value of such reformulation is unlocked when one computes (approximate) hypergradients via appropriate continuous relaxation or discrete gradient estimators [niculae2023discrete]. These techniques may be combined with the procedures described in this chapter, making it possible to avoid the exponential cost of computing $f$ on all the configuration. This is still an active area of research: successful examples include the optimization of graph structures in transductive learning [franceschi2019learning] and topological order of nodes in DAG learning [zantedeschi2023dag], deep learning architectures [liu2018darts] and dropout coefficients [gal2017concrete].
@@ 7.7 Discussion
Implicit differentiation methods are “oblivious” to the way the approximate solution $w_{T}$ has been found. This means that they have a computational advantage in memory over the reverse-mode differentiation as they do not require storing the entire dynamics. If implemented using algorithmic differentiation, they have a run-time advantage over forward-mode differentiation since they only deal with vector quantities (the $q_{T,K}$’s). However, for this very reason, they do not directly allow computing the gradient of hyperparameters related to the optimization of the inner problem, such as the initialization mapping or the learning rate. Furthermore, they are also quite sensitive to the proprieties of the inner problem (e.g. for the fixed-point method, the dynamics must be a contraction), and therefore are applicable to a narrower class of problems. A comparative theoretical study between the three main approaches covered in this chapter is presented in [grazzi2020iteration].
Computational disadvantages of plain iterative reverse mode are related to memory requirements. Indeed, the memory complexity is $O(TW)$, with $W$ the number of weights, which can quickly become prohibitive when tuning the hyperparameters associated with modern networks. The problem has been addressed in different ways. maclaurin2015gradient suggested to completely avoid the storage of $s_{t}$, recomputing it in the backward pass by reversing the learning dynamics (7.9). Since the operation is numerically unstable, reconstructed initial weights may be very different from the original ones. As a remedy, they proposed an effective numerical strategy capable of overcoming the loss of information due to finite precision arithmetic. A general technique for reducing the memory requirements in reverse mode AD is checkpointing [Griewank08:_evaluat_deriv, Ch. 12], where only a subset of the possible checkpoints is stored, namely $w_{jC}$ for $j=1,\dots\lceil T/C\rceil$, being $C$ a given window size. In so doing, one must redo the forward pass from $jC$ to $(j+1)C-1$ to recover the missing $w_{t}$, effectively doubling the computational effort. It is easy to see that the best window size is $O(\sqrt{T})$, which allows us to complete hypergradient calculations with a memory size of $O(W\sqrt{T})$. This technique has been considered in [shaban2019truncated], together with a more radical truncation strategy similar to truncated backpropagation through time [DBLP:journals/neco/WilliamsP90], which however introduces further bias in the computation of hypergradients.
@@
In the last five chapters we covered the main classes of algorithms for hyperparamter optimization. Here, moving away from particular techniques, we discuss horizontally advanced topics in HPO such as multi-objective and constrained optimization (Section 8.1), online HPO and hyperparameter schedules (Section 8.2) and neural architecture search (Section 8.3). Next, we dedicate Section 8.4 to meta-learning, a closely related topic in the literature, discussing similarities and differences with the typical hyperparameter optimization setting. We conclude the chapter with a brief discussion on transfer of hyperparamters across model sizes, a novel topic that is relevant especially for the tuning of large foundational models (Section 8.5).

### v3

@@ 7.3 Hypergradients via Iterative Differentiation
Another approach to computing hypergradients involves explicitly accounting for an iterative training dynamics. Methods based on iterative differentiation further assume that solutions to the learning algorithm problem of Eq. (7.3) are sought (with approximation) by iterating an optimization dynamics for $T$ steps:
@@
We call $s_{t}=[w^{\mathsf{T}}_{t},v^{\mathsf{T}}_{t}]^{\mathsf{T}}\in\mathbb{R}^{d}$ the “state” of the optimizer, which contains the current parameters, $w_{t}$, and a (possibly empty) set of auxiliary variables $v_{t}$ such as velocities and moments for algorithms like Adam and RMSProp [ruder:2017:overview_gradient_descent]. This process yields a solution $w_{T}=w_{T}(\lambda)$. The transition maps $\Phi_{t}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{d}\times\mathbb{R}^{m}\times\mathscr{D}\rightarrow\mathbb{R}^{d}$ are assumed to be differentiable and describe the actual operations associated with a specific optimizer. Typically, $\Phi_{t}$ depends explicitly on the current minibatch $\mathcal{D}_{t}$ (e.g., for stochastic gradient descent) and it internally requires the computation of the inner objective gradients:
@@
Note that $\Phi$ depends on $\lambda$ both explicitly and implicitly via $s_{t-1}$. Also, $\lambda$ might contain quantities related to the optimizer itself, such as the learning rate. As an example, the dynamics (7.9) for stochastic gradient descent with momentum are
@@
In a nutshell, iterative differentiation schemes for gradient-based HPO rely on the application of automatic differentiation (AD) to the composite map $\mathscr{M}\circ\Phi_{T}\circ\dots\circ\Phi_{1}$. We briefly survey AD next, and refer the reader to the book by Griewank08:_evaluat_deriv for an in-depth account.
@@ 7.3.1 Automatic Differentiation
Automatic (or algorithmic) differentiation (AD) [Griewank08:_evaluat_deriv, baydin2017automatic] is the standard tool for computing gradients in deep learning. Suppose we have a generic differentiable function $G\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{N}\mapsto\mathbb{R}^{M}$ with $(\eta_{1},\dots,\eta_{M})=G(\xi_{1},\dots,\xi_{M})$ and we are interested in computing the Jacobian matrix, with entries $\frac{\partial\eta_{i}}{\partial\xi_{j}}$. In AD, we begin by describing $G$ as a computational graph $(V,E)$, which is a directed acyclic graph whose nodes $V$ include inputs $\xi_{j}$, outputs $\eta_{i}$, and intermediate variables that are computed by applying elementary functions to inputs or other intermediate variables, that can be scalars, vectors or tensors. Elementary functions are typically atomic operations such as (tensor) sum, (Hadamard) product, and (elementwise) transcendental functions, whose partial derivatives are readily known. Edges $E$ in the graph describe functional dependencies (i.e., they link intermediate variables to the elementary functions they are fed into). Forward propagation, following a topological sort of the graph, allows us to compute $G(\xi_{1},\dots,\xi_{M})$. Two AD strategies are available, forward and reverse mode, which are based on the definition of local derivatives and a message-passing procedure based on the standard chain rule of calculus.
@@
initialized as $\dot{\xi_{j}}=1$, where $\pi_{k}$ are $v_{k}$’s parents, and iterated following a topological sort of $(V,E)$. In the above equation, if $v_{k}$ and $v_{\ell}$ are vector-valued, $\frac{\partial v_{k}}{\partial v_{\ell}}$ denotes a Jacobian matrix. The desired derivatives can be finally read as $\dot{\eta}_{i}$. The same propagation allows us to compute the $j$-th column of the Jacobian and to obtain the whole Jacobian, where the propagation needs to be repeated for each input $\xi_{j}$. Assuming elementary functions can be computed in $O(1)$, the total computational cost is $O(N|E|)$. Forward mode is thus advantageous for functions with few inputs and many outputs. If on the other hand $N=O(|E|)$, as it happens for example for the objective function of neural networks, forward mode is as inefficient as computing derivatives numerically. It is therefore not surprising that forward mode is never used to train neural network models (although there is a remarkable exception when training recurrent networks on data streams [williams_learning_1989], a setting where reverse mode cannot be applied).
In reverse mode, the strategy used by the well known backpropagation algorithm [rumelhart1986learning], the local derivative of $v_{k}$, usually called adjoint, is defined as
@@ 7.3.2 Computing Hypergradients
We are now ready to apply the principles of AD to the differentiation of the validation error with respect to $\lambda$. A simplified view of the computational graph mapping $\lambda$ into the response $f(\lambda)$ is shown in Figure 7.1, where minibatches $\mathcal{D}_{t}$ and initial weights $w_{0}$ are supplied as given constants [^1] Note that $w_{0}$ could be equally well included in $\lambda$ but here we keep it distinct for clarity.. Here the number of inputs $M$ is the dimension of the hyperparameter vector and there is a scalar ($N=1$) output. Forward and reverse calculations are presented in the work of franceschi2017forward and repeated below.
@@
$\nabla f(\lambda)=\left.\frac{\partial\mathcal{M}(u,\mathcal{V})}{\partial u}\right|_{u=w_{T}(\lambda)}\dot{s}_{T}$
@@
$\bar{s}_{T}=\left.\frac{\partial\mathcal{M}(u,\mathcal{V})}{\partial u}\right|_{u=w_{T}(\lambda)}$
@@ 7.4 Fixed-Point Methods
Another possible approach that sits in between the iterative and the implicit differentiation schema is given by the so-called fixed-point method. The method is based on the observation that one can rewrite the condition of Eq. (7.4) as a fixed point equation of an appropriate contractive dynamics
@@
which could be, for instance, gradient descent $\Phi(w,\lambda)=w-\eta\nabla\mathcal{J}$.  [^2] As in the iterative case, the dynamics may be defined on an augmented state, for example when using gradient descent with momentum. Then, the application of the implicit function theorem to Eq. (7.15) yields to
@@
and, repeating the procedure described above, Eq. (7.7) becomes
@@
Starting from any point $q_{T,0}\in\mathbb{R}^{d}$, one can iterate
@@
to find an approximate fixed-point of Eq. (7.16) and then plug $q_{T,K}$ into Eq. (7.8) to obtain a hypergradient.
This computational scheme emerges in more general cases where $\Phi$ is not necessarily an optimization dynamics: $\Phi$ may represent a stable recurrent neural network [miller2019stable], a graph neural network [scarselli2009graph] or an equilibrium model [bai2019deep]. In these contexts, the fixed-point algorithm is usually known as recurrent backpropagation [almeida1987learning, liao2018reviving]. Warm-starting $q_{T,0}$ with the vector obtained at the previous optimization iteration (of $\lambda$) may speed up the convergence of Eq. (7.17).
@@ 7.5 Use Cases
One possibility is to attach a hyperparameter to every weight in a deep network. We have already mentioned the idea proposed by maclaurin2015gradient regarding the initial values for the weights as hyperparameters. In the same paper, the authors also suggest a closely related opportunity where a regularization penalty is associated to every individual weights. The meta-learning technique presented by franceschi2018bilevel can be interpreted as treating all weights in the first layers of a deep network as hyperparameters, and optimize the loss on new tasks with respect to the weights in the classification head.
Another direction that has been studied is to attach hyperparameters to data. maclaurin2015gradient demonstrated an extreme experiments where image pixels themselves are hyperparameters. In a more realistic scenario, hypergradients have been used to tune the hyperparameters of differentiable data augmentation procedures, yielding significant speedups without performance loss compared to other data augmentation searching methods [lin:2019:online_hyperparameter_learning, hataya:2020:faster_autoaugment_learning]. A different opportunity is to assign a cost weight to each example in a supervised learning setting so that the overall loss for a training set of $n$ examples is $\sum_{i=1}^{n}\lambda^{(i)}\mathcal{J}(h(x^{(i)}),y^{(i)})$. Here we have $n$ hyperparameters $\lambda^{(i)}$. franceschi2017forward show how this idea can be exploited to filter out from the training set examples with noisy (or perhaps poisoned) labels. In this regard, it is interesting to note that (although not directly related to HPO) the problem of building training set attacks on linear models can be formulated in terms of bilevel optimization [mei:2015:using_machine_teaching] and has commonalities with gradient-based HPO in its mathematical structure.
@@
7.5.1 Probabilistic Reparameterization of Discrete Hyperparameters
The presence of discrete hyperparameters in the response function — and therefore in the training objective or dynamics — prevents a straightforward application of gradient-based HPO techniques we just described. For some of such hyperparameters, however, it might make sense to reformulate them as random variable that follow some discrete distribution. For instance, consider a discrete binary choice $\lambda\in\{0,1\}$ (e.g., indicating whether to use an $L_{1}$ or $L_{2}$ regularizer in Eq. (2.3)). Then, one may define a discrete binary random variable $\iota$ as follows:
@@ 7.5.1 Probabilistic Reparameterization of Discrete Hyperparameters
and reinterpret $\lambda$ as a drawing of $\iota$, where $\mathrm{Ber}(\theta)$ is the Bernoulli distribution with success parameter $\theta$. Similarly, one may reformulate hyperparameters encoding multiple choices via categorical distributions. Crucially, the distribution parameter $\theta$ is now continuous and one may consider optimizing the response function
@@
instead of $f$, having that $\min_{\theta}g(\theta)=\min_{\lambda}f(\lambda)$. Although Eq. (7.19) might look like a circular argument, the value of such reformulation is unlocked when one computes (approximate) hypergradients via appropriate continuous relaxation or discrete gradient estimators [niculae2023discrete]. These techniques may be combined with the procedures described in this chapter, making it possible to avoid the exponential cost of computing $f$ on all configurations. This is still an active area of research: successful examples include the optimization of graph structures in transductive learning [franceschi2019learning] and topological order of nodes in DAG learning [zantedeschi2023dag], deep learning architectures [liu2018darts] and dropout coefficients [gal2017concrete].
@@ 7.6 Discussion
Implicit differentiation methods are “oblivious” to the way the approximate solution $w_{T}$ has been found. This means that they have a computational advantage in memory over the reverse-mode differentiation as they do not require storing the entire dynamics. If implemented using algorithmic differentiation, they have a run-time advantage over forward-mode differentiation since they only deal with vector quantities (the $q_{T,K}$’s). However, for this very reason, they do not directly allow computing the gradient of hyperparameters related to the optimization of the inner problem, such as the initialization mapping or the learning rate. Furthermore, they are also quite sensitive to the proprieties of the inner problem (e.g., for the fixed-point method, the dynamics must be a contraction), and therefore are applicable to a narrower class of problems. A comparative theoretical study between the three main approaches covered in this chapter is presented in the work by grazzi2020iteration.
Computational disadvantages of plain iterative reverse mode are related to memory requirements. Indeed, the memory complexity is $O(TW)$, with $W$ the number of weights, which can quickly become prohibitive when tuning the hyperparameters associated with modern networks. The problem has been addressed in different ways. maclaurin2015gradient suggested to completely avoid the storage of $s_{t}$, recomputing it in the backward pass by reversing the learning dynamics of Eq. (7.9). Since the operation is numerically unstable, reconstructed initial weights may be very different from the original ones. As a remedy, they proposed an effective numerical strategy capable of overcoming the loss of information due to finite precision arithmetic. A general technique for reducing the memory requirements in reverse mode AD is checkpointing [Griewank08:_evaluat_deriv, Ch. 12], where only a subset of the possible checkpoints is stored, namely $w_{jC}$ for $j=1,\dots\lceil T/C\rceil$, being $C$ a given window size. In so doing, one must redo the forward pass from $jC$ to (j+1)C1(j+1)C-1 to recover the missing $w_{t}$, effectively doubling the computational effort. It is easy to see that the best window size is $O(\sqrt{T})$, which allows us to complete hypergradient calculations with a memory size of $O(W\sqrt{T})$. This technique has been considered by shaban2019truncated, together with a more radical truncation strategy similar to truncated backpropagation through time [DBLP:journals/neco/WilliamsP90], which however introduces further bias in the computation of hypergradients.
@@
In the last five chapters we covered the main classes of algorithms for hyperparameter optimization. Here, moving away from particular techniques, we discuss horizontally advanced topics in HPO such as multi-objective and constrained optimization (Section 8.1), online HPO and hyperparameter schedules (Section 8.2) and neural architecture search (Section 8.3). Next, we dedicate Section 8.4 to meta-learning, a closely related topic in the literature, discussing similarities and differences with the typical hyperparameter optimization setting. We conclude the chapter with a brief discussion on transfer of hyperparameters across model sizes, a recent topic that is relevant especially for the tuning of large foundational models (Section 8.5).